# GeoGuess: Multimodal Reasoning based on Hierarchy of Visual Information in Street View

### v1

@@
Abstract.
Multimodal reasoning is a process of understanding, integrating and inferring information across different data modalities. It has recently attracted surging academic attention as a benchmark for Artificial Intelligence (AI). Although there are various tasks for evaluating multimodal reasoning ability, they still have limitations. Lack of reasoning on hierarchical visual clues at different levels of granularity, e.g., local details and global context, is of little discussion, despite its frequent involvement in real scenarios. To bridge the gap, we introduce a novel and challenging task for multimodal reasoning, namely GeoGuess. Given a street view image, the task is to identify its location and provide a detailed explanation. A system that succeeds in GeoGuess should be able to detect tiny visual clues, perceive the broader landscape, and associate with vast geographic knowledge. Therefore, GeoGuess would require the ability to reason between hierarchical visual information and geographic knowledge. In this work, we establish a benchmark for GeoGuess by introducing a specially curated dataset GeoExplain which consists of panoramas-geocoordinates-explanation tuples. Additionally, we present a multimodal and multilevel reasoning method, namely SightSense which can make prediction and generate comprehensive explanation based on hierarchy of visual information and external knowledge. Our analysis and experiments demonstrate their outstanding performance in GeoGuess.
Multimodal Reasoning, Multimodal Dataset, Image Geo-Localization, Hierarchical Visual Information
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generation
Figure 1. An GeoExplain example: panoramas-geoordinates-explanation tuple.
1. Introduction
@@
Humans can not only see the world, but more importantly comprehend the world through it. Observing, associating visual or textual knowledge, making interpretation, and even building a cognitive framework demonstrate a critical and widely used ability of human intelligence, known as multimodal reasoning. For example, a human can roughly infer the meaning of a traffic sign based on disassembled components and knowledge even if they have never seen it before. Therefore, developing more human-like multimodal reasoning ability is an essential step towards advancing Artificial General Intelligence (AGI) (Goertzel and Pennachin, 2007).
@@
In recent years, there has been significant enthusiasm for researching on Large Language Model (LLM) (Radford et al., 2018; Brown et al., 2020; Devlin et al., 2019; Touvron et al., 2023) driven by the rise of self-attention mechanism and Transformer architecture (Vaswani et al., 2017). We have witnessed the remarkable performance of LLM on Natrual Language Understanding across various Natrual Language Processing (NLP) generative tasks (Min et al., 2023; Raffel et al., 2020; Yang et al., 2019). The great progress prompted the AI community to embark on an exploration of Multimodal Large Language Model (MLLM) (Radford et al., 2021; Alayrac et al., 2022; Li et al., 2023) and multimodal reasoning (Antol et al., 2015; Wei et al., 2024) that incorporate visual context as additional input. Despite the success achieved by techniques such as the Chain-of-Thought (CoT) (Wei et al., 2022), Tree-of-Thoughts (ToT) (Yao et al., 2023), and Self-Consistency (Wang et al., 2022), some studies highlighted that state-of-the-art MLLMs still struggle in multimodal reasoning (Fu et al., 2024; Samadarshi et al., 2024).
@@
In response, a diverse range of multimodal reasoning testbeds have been developed, such as Science Question Answering (ScienceQA) (Lu et al., 2022), Geometric Question Answering (Chen et al., 2021) and GQA (Hudson and Manning, 2019). Although these testbeds vary in domains and disciplines, we believe all of them still face the following limitations:
@@
  • Superficial Reasoning Process: These testbeds were relatively simple, requiring only superficial reasoning process and limited external knowledge which cannot embody the diverse, hierarchical and complex nature of human reasoning in real-world situations. This limitation weakens the practical grounding of expert-level multimodal reasoning in real-world settings.
@@
  • Insufficient Visual Context at Different Granularity: Existing multimodal reasoning testbeds primarily focus on reasoning over global visual context, often neglecting the critical role of local visual clues throughout the whole reasoning process. Identifying tiny visual clues and extracting meaningful information from them is also an intuitive aspect of human cognition. Lack of multi-level visual understanding restricts reasoning performance on detail-oriented tasks.
@@
To this end, we introduce a new task for multimodal reasoning: GeoGuess. Given a street view image, the task is to predict the location depicted in the image and provide a comprehensive explanation. GeoGuess is a variant of image geo-localization task (Hays and Efros, 2008; Zhou et al., 2024). Existing exploration of image geo-localization only requires predicting coordinates of the location in a black box manner. On the contrary, GeoGuess places greater emphasis on reasonable interpretations in natural language as illustrated in Figure 1. GeoGuess is one of the most complex and challenging tasks designed for multimodal reasoning. Participants must analyze a variety of visual elements, including global context like landscape and specific local details like car plate, while also integrating external knowledge. The complexity reflects the multifaceted nature of real-world reasoning, making it a challenging endeavor even for humans. Therefore, a system that excels in GeoGuess must identify tiny clues in the image, grasp the overall vibe, associate geographic knowledge, reason between these elements and draw a final inference.
@@
Furthermore, we introduce a novel dataset, GeoExplain, collected from the image geo-localization game GeoGuessr (GeoGUessr, 2025) in which players must deduce locations from Google Street View imagery. As Figure 1 shows, each sample in GeoExplain includes a set of panoramic images, the geo-coordinates of the location, a text description of the location and an explanation of how the exact location can be inferred from the image contents. We remove locations that are easy to recognize to make GeoGuess an expert-level task even for humans. Building on the strengths of the novel dataset, we also propose a three-stage method, SightSense, specially designed for hierarchical multimodal reasoning to leverage local and global visual context and external knowledge. More specifically, in the first stage, the model identifies fine-grained visual clues via an open-vocabulary object detection. The detected objects are then used to retrieve knowledge snippets in the second stage. Finally, these elements are fused through an adapter, enabling the model to generate a comprehensive response. The effectiveness and efficiency of SightSense has been demonstrated through our experiments on GeoExplain, outperforming state-of-the-art methods.
@@
2. Related Works
2.1. Image Geo-Localization
@@
Recent years have witnessed remarkable advancements in the field of image geo-localization, a subdomain of computer vision. The task of image geo-localization involves estimating the precise geographic coordinates where the photograph was taken. Recent developments in image geo-localization can be categorized into two main approaches: retrieval-based method, classification-based method and LLM-based method.
@@
The retrieval-based method (Regmi and Shah, 2019; Shi et al., 2019, 2020; Toker et al., 2021; Zhu et al., 2021) searches for the most similar photo within a large-scale photo set in which each photo is annotated with its exact geographical coordinates. The coordinates of the best-matched image are then identified as the result. The classification-based method (Muller-Budack et al., 2018; Pramanick et al., 2022; Seo et al., 2018; Vo et al., 2017; Weyand et al., 2016) approaches the geo-localization task as an earth segments classification problem, employing deep learning networks in a black-box manner. Recent LLM-based method (Liu et al., 2024; Han et al., 2024; Yan and Lee, 2024; Li et al., 2024) utilize the reasoning ability obtained from large-scale training corpus to generate prediction in natural language.
@@
Despite the significant success achieved by these methods, each approach still faces notable challenges. For retrieval-based method, matching images captured from different viewpoints at the same location remains difficult, and the approach demands substantial computational resource due to searching in an additional large database. Besides, some local clues may be overlooked because they may be disappeared in different photos even at the same location. Regression-based method suffers from lack of interpretability. As the regression process functions in a black-box manner, it is difficult to understand how predictions are made, making the results less transparent and less persuasive. Neither approach follows the human way of thinking, which involves sensing the global vibe from a photo, such as terrain or landscape, identifying subtle details, such as plants or animals, and connecting them with geographical knowledge to infer the exact location. On the other hand, existing LLM-based method suffers from poor performance on non-landmark scene and hallucination.
2.2. Visual Language Modelling
@@
Scaling up the size of models’ parameters and training data has significantly enhanced the capabilities of attention-based Large Language Models (LLMs) (Radford et al., 2018; Brown et al., 2020; Devlin et al., 2019), enabling them to surpass human performance in various generative tasks in Natural Language Processing (NLP), such as question answering (Rabinovich et al., 2023; Chen et al., 2022) and text summarization (Liu et al., 2024a; Zhang et al., 2020). However, most of them were limited by incapacity of reading visual information. Conversely, Large Vision Model (LVMs) (Dosovitskiy et al., 2020; Liu et al., 2021; Guo et al., 2022) can see clearly but cannot talk. Building on the success of both LLM and LVM, significant research efforts have been dedicated to extending the impressive zero-shot capabilities of LLMs to the visual domain by incorporating visual modalities of LVM into generative tasks, which is called Multimodal Large Language Model (MLLM), such as GPT-4 (Achiam et al., 2023), LLaVA (Liu et al., 2023) and Llama (Grattafiori et al., 2024). An MLLM can bridge effective connections between different gaps, thus succeeding in various multimodal tasks, including 1) image-grounded question-answering dialogue system (Das et al., 2017; Murahari et al., 2020; Nguyen et al., 2019), 2) image-grounded chit-chat dialogue system (Shuster et al., 2020, 2021), 3) visual-evidence-embedded dialogue system (Shen et al., 2021; Liang et al., 2021), 4) image-response dialogue system (Zang et al., 2021; Sun et al., 2021), 5) video-grounded question-answering dialogue system (Le et al., 2019; Alamri et al., 2019; Le et al., 2021), 6) video-grounded chit-chat dialogue system (Han et al., 2023; Lin et al., 2023; Cheng et al., 2024) and 7) video-grounded real-time comment dialogue (Ma et al., 2019; Wang et al., 2020).
2.3. Knowledge-Intensive Visual Reasoning
Except bridging modality gaps and enabling multi-granularity, reducing hallucination remains an evident challenge in developing convincing MLLMs (Huang et al., 2025). Most notably issues related to hallucination happen to knowledge-intensive visual language tasks, e.g., knowledge-intensive VQA (Marino et al., 2019; Schwenk et al., 2022), multi-hop VQA (Chang et al., 2022) and visual reasoning (Chung et al., 2024; Johnson et al., 2017).
Knowledge-intensive VQA requires various external expert knowledge which may not be fully covered by a MLLM’s internal parametric knowledge to support accurate answer generation. To mitigate hallucination caused by the absence of expert knowledge, Retrieval Augmented Generation (RAG) (Lewis et al., 2020) leverage a search engine for external knowledge retrieval, assisting answer generation (Gao et al., 2023). Visual reasoning aims to equip machines with the ability to analyze visual information and make decisions or inferences based on that. Studies (Fu et al., 2024; Samadarshi et al., 2024) have demonstrated visual reasoning abilities in MLLM still falls behind human performance in a variety of tasks, despite the emergence of Chain-of-Thought (CoT) prompting and related techniques (Wei et al., 2022; Yao et al., 2023).
3. GeoExplain
@@
(1) $G=\{(\mathbf{P_{i}},L_{i},E_{i})\}_{i=1}^{n}$
Each tuple corresponds to a specific geographic location and $n$ is the total number of samples. The panorama $\mathbf{P_{i}}=\{I_{1},I_{2},\dots,I_{im}\}$ is a series of street view images of the location from different angles, $L_{i}$ is the exact location on street level, and the explanation $E_{i}$ provides the expert-level natural-language description how human can infer the location from the visual information in $\mathbf{P_{i}}$. Figure 1 demonstrates an example of GeoExplain. We collected GeoExplain from an image geo-localization game GeoGuessr and its community on Reddit (Reddit, 2025).
We also collected a knowledge base, GeoKnowledge, which consists of geographical knowledge to assist reasoning. GeoKnowledge contains image-snippets pairs indexed by country where the snippets tell how to identify a country from visual clues in the image. GeoKnowledge can be represented by Equation 2, where $KI_{j}$ is an image of country-specific object and $S_{j}$ is the corresponding knowledge snippets.
(2) $K=\{(KI_{j},S_{j})\}_{j=1}^{k}$
@@
3.1. Data Collection
@@
After carefully filtering locations, we collected street view images in different angles from Google Street Views (Google, 2025). On the other hand, we extracted corresponding expert-level reasoning explanations from Reddit discussions. Specifically, posters usually include their scores in the discussion, so we first splited the game scores and reasoning processes. Then, we filtered these explanations as follows: 1) Remove all explanations with scores lower than 4000 (with max 5000), 2) Remove all explanations that are not in English, 3) Remove all explanations that rely on non-reasoning factors to arrive at the answer, such as living nearby or having previously finished the same challenge.
We collected GeoKnowledge also from a geographical knowledge website which is established specially for sloving the image-geolocalization puzzles.
3.2. Data Analysis
We present an analysis of GeoExplain in this section. There are in total 5459 locations, 169972 panorama images and 41068 explanations on how to find the location based on images. We endeavor to ensure the richness and diversity of vocabulary in these explanations. The locations in GeoExplain are distributed across all continents, covering a wide range of terrains as shown in Table 1. They include diverse landscapes such as mountains, deserts, forests, coastlines, and urban environments, ensuring a comprehensive geographical representation. Regarding panorama statistics, each location contains an average of 31.13 street view image snippets, with a maximum 33 and a minimum 3. Their size vary, ranging from (600 * 280) to (512 * 512).
AF AN AS EU NA OC SA
445 3 1434 1730 771 175 992
Table 1. Location Distribution across All Continents
@@ 3.2. Data Analysis
Figure 2. Human Analysis on GeoExplain and Other Multimodal Reasoning Dataset
@@
4. SightSense
Figure 3. Architecture of SightSense: Visual Clue Detection, Multimodal Knowledge Retrieval and Reasoning Generation.
@@ 4.1. Problem Statement
(3) $L^{*},E^{*}=\arg\max_{L,E}P(L,E|P,K;\theta)$
4.2. Model Architecture
In order to fully leverage visual information at different levels of granularity, SightSense uses a three-stage method. Figure 3 illustrates the overall workflow. The three stages are Visual Clues Detection, Multimodal Knowledge Retrieval and Reasoning Generation respectively. In the first stage, potentially identifiable local elements are extracted from street view images $\mathbf{P}$ by using an open-vocabulary object detector. The second stage retrieves relevant geographical knowledge $KI,S$ based on these visual clues. Finally, the reasoning stage integrates global context, detected visual clues and retrieved external knowledge to generate informative answer and explanation ${L,E}$.
4.3. Visual Clue Detection

### v2

@@
Abstract
Multimodal reasoning is a process of understanding, integrating and inferring information across different data modalities. It has recently attracted surging academic attention as a benchmark for Artificial Intelligence (AI). Although there are various tasks for evaluating multimodal reasoning ability, they still have limitations. Reasoning on hierarchical visual clues at different levels of granularity, i.e., local details and global context, is of little discussion, despite its frequent involvement in human reasoning. To bridge the gap, we introduce a novel and challenging task for multimodal reasoning, namely GeoGuess. Given a street view image, the task is to identify its location and provide a detailed explanation. In this work, we establish a benchmark for GeoGuess by introducing a specially curated dataset GeoExplain which consists of panoramas-geocoordinates-explanation tuples. Additionally, we present a multimodal and multilevel reasoning method, namely SightSense which can make prediction and generate comprehensive explanation. Our analysis and experiments demonstrate their outstanding performance in GeoGuess.
1 Introduction
@@
Figure 1: An GeoExplain example: panoramas-geoordinates-explanation tuple. Each sample consists of a set of panoramas, geoordinates and explanation on how to infer the location. Hierarchical visual information, i.e., local details and global context are both necessary in this example.
@@
Humans can not only see the world, but more importantly comprehend the world through it. Observing, associating visual or textual knowledge, making interpretation, and even building a cognitive framework demonstrate a critical and widely used ability of human intelligence, known as multimodal reasoning. Therefore, developing more human-like multimodal reasoning ability is an essential step towards advancing Artificial General Intelligence (AGI) [15].
@@
In recent years, there has been significant enthusiasm for researching on Large Language Model (LLM) [49, 6, 11, 63] driven by the rise of self-attention mechanism and Transformer architecture [64]. We have witnessed the remarkable performance of LLM on Natrual Language Understanding [43, 51, 76]. The great progress prompted the AI community to embark on an exploration of Multimodal Large Language Model (MLLM) [50, 3, 30] and multimodal reasoning [4, 72] that incorporate visual context as additional input. Despite the success achieved by techniques such as the Chain-of-Thought (CoT) [71], Tree-of-Thoughts (ToT) [77], and Self-Consistency [70], some studies highlighted that state-of-the-art MLLMs still struggle in multimodal reasoning [13, 54].
@@
In response, a diverse range of multimodal reasoning testbeds have been developed, such as Science Question Answering (ScienceQA) [41], Geometric Question Answering [7] and GQA [24]. Although these testbeds vary in domains and disciplines, we believe they still face the following limitations:
@@
  • Superficial Reasoning Process: These testbeds were relatively simple, requiring only superficial reasoning process and limited external knowledge which cannot embody the diverse, hierarchical and complex nature of human reasoning. This limitation weakens the practical grounding of expert-level multimodal reasoning.
@@
  • Insufficient Visual Context at Different Granularity: Existing multimodal reasoning testbeds primarily focus on reasoning over global visual context, often neglecting local visual clues. Identifying tiny visual clues and extracting meaningful information from them is also an intuitive aspect of human cognition. Lack of multi-level visual understanding restricts reasoning performance on detail-oriented tasks.
@@
To this end, we introduce a new task for multimodal reasoning: GeoGuess. Given a street view image, the task is to predict the location depicted in the image and provide a comprehensive explanation. GeoGuess is a variant of image geo-localization task [22, 81]. Existing exploration of image geo-localization only requires predicting coordinates of the location in a black box manner. On the contrary, GeoGuess places greater emphasis on reasonable interpretations in natural language as illustrated in Figure 1. GeoGuess is one of the most complex and challenging tasks designed for multimodal reasoning. Participants must analyze hierarchical visual elements, including global context like landscape and local details like car plate. The complexity reflects the multifaceted nature of real-world reasoning, making it a challenging endeavor even for humans. Therefore, a system that excels in GeoGuess must identify tiny clues in the image, grasp the overall vibe, associate geographic knowledge, reason between these elements and draw a final inference.
@@
Furthermore, we introduce a novel dataset, GeoExplain, collected from the image geo-localization game GeoGuessr [14]. Each sample in GeoExplain includes a set of panoramic images, the geo-coordinates of the location and an explanation of how the exact location can be inferred from the image contents. We remove locations that are easy to recognize to make GeoGuess an expert-level task even for humans. We also propose a three-stage method, SightSense, specially designed for hierarchical multimodal reasoning to leverage local and global visual understanding and external knowledge. More specifically, in the first stage, the model identifies fine-grained visual clues via an open-vocabulary object detection. The detected objects are then used to retrieve knowledge snippets in the second stage. Finally, these elements and global context are fused through an adapter, enabling the model to generate a comprehensive response. Our experiments demonstrates the effectiveness of SightSense on GeoGuess, outperforming other state-of-the-art baselines.
@@
2 Related Works
2.1 Image Geo-Localization
@@
Recent years have witnessed remarkable advancements in the field of image geo-localization, a subdomain of computer vision. The task of image geo-localization involves estimating the precise geographic coordinates where the photograph was taken. Recent developments in image geo-localization can be categorized into three main approaches: retrieval-based method, regression-based method and LLM-based method.
@@
The retrieval-based method [53, 57, 58, 62, 82] searches for the most similar photo within a large-scale photo set in which each photo is annotated with its exact geographical coordinates. The coordinates of the best-matched image are then identified as the result. The regression-based method [19, 66, 47, 55, 67, 73] approaches the geo-localization task as an earth segments regression problem, employing deep learning networks in a black-box manner. Recent LLM-based method [38, 21, 75, 31] utilize the reasoning ability obtained from large-scale training corpus to generate prediction in natural language.
@@
Despite the significant success achieved by these methods, each approach still faces notable challenges. For retrieval-based method, matching images captured from different viewpoints at the same location remains difficult, and the approach demands substantial computational resource due to searching in an additional large database. Besides, some local clues may be overlooked because they may be disappeared in different photos even at the same location. Regression-based method suffers from lack of interpretability. As the regression process functions in a black-box manner, it is difficult to understand how predictions are made, making the results less transparent and less persuasive. Neither approach follows the human way of thinking, which involves sensing the global vibe from a photo, such as terrain or landscape, identifying subtle details, such as plants or animals, and connecting them with geographical knowledge to infer the exact location. On the other hand, existing LLM-based method suffers from poor performance on non-landmark scene and hallucination [23].
2.2 Visual Language Modelling
@@
Scaling up the size of models’ parameters and training data has significantly enhanced the capabilities of attention-based Large Language Models (LLMs) [49, 6, 11], enabling them to surpass human performance in various generative tasks in Natural Language Processing (NLP), such as question answering [48, 8] and text summarization [39, 79]. However, most of them were limited by incapacity of reading visual information. Conversely, Large Vision Model (LVMs) [12, 40, 18] can see clearly but cannot talk. Building on the success of both LLM and LVM, significant research efforts have been dedicated to extending the impressive zero-shot capabilities of LLMs to the visual domain by incorporating visual modalities of LVM into generative tasks, which is called Multimodal Large Language Model (MLLM), such as GPT-4 [1], LLaVA [35] and Llama [17]. An MLLM can bridge effective connections between different gaps, thus succeeding in various multimodal tasks, including 1) image-grounded question-answering dialogue system [10, 44, 45], 2) image-grounded chit-chat dialogue system [59, 60], 3) visual-evidence-embedded dialogue system [56, 32], 4) image-response dialogue system [78, 61], 5) video-grounded question-answering dialogue system [26, 2, 27], 6) video-grounded chit-chat dialogue system [20, 34, 9] and 7) video-grounded real-time comment dialogue [42, 69].
3 GeoExplain
@@
$G=\{(\mathbf{P_{i}},L_{i},E_{i})\}_{i=1}^{n}$ (1)
Each tuple corresponds to a specific geographic location and $n$ is the total number of samples. The panorama $\mathbf{P_{i}}=\{I_{1},I_{2},\dots,I_{im}\}$ is a series of street view images of the location from different angles, $L_{i}$ is the exact location on street level, and the explanation $E_{i}$ provides the expert-level natural-language description how human can infer the location from the visual information in $\mathbf{P_{i}}$. Figure 1 demonstrates an example of GeoExplain. We collected GeoExplain from an image geo-localization game GeoGuessr and its community on Reddit [52].
@@
3.1 GeoKnowledge
To incorporate external geographic knowledge, we collect a domain-specific knowledge base, GeoKnowledge. GeoKnowledge consists of image-snippets pairs indexed by country where the snippets tell how to identify a country from visual clues in the image. GeoKnowledge can be represented by Equation 2, where $KI_{j}$ is an image of country-specific object and $S_{j}$ is the corresponding knowledge snippets.
$K=\{(KI_{j},S_{j})\}_{j=1}^{k}$ (2)
3.2 Data Collection
@@
After carefully filtering locations, we collected street view images in different angles from Google Street Views [16]. On the other hand, we extracted corresponding expert-level reasoning explanations from Reddit discussions. Specifically, posters usually include their scores in the discussion, so we first splited the game scores and reasoning processes. Then, we filtered these explanations as follows: 1) Remove all explanations with scores lower than 4000 (with max 5000), 2) Remove all explanations that are not in English, 3) Remove all explanations that rely on non-reasoning factors to arrive at the answer, such as living nearby or having previously finished the same challenge.
GeoKnowledge is collected from a geographical knowledge website [^1] Plonk It which is established specially for sloving the image-geolocalization puzzles.
3.3 Data Analysis
We present an analysis of GeoExplain in this section. There are in total 5459 locations, 169972 panorama images and 41068 explanations on how to find the location based on images. We endeavor to ensure the richness and diversity of vocabulary in these explanations. The locations in GeoExplain are distributed across all continents, covering a wide range of terrains. They include diverse landscapes such as mountains, deserts, forests, coastlines, and urban environments, ensuring a comprehensive geographical representation. Regarding panorama statistics, each location contains an average of 31.13 street view image snippets, with a maximum 33 and a minimum 3. Their size vary, ranging from (600 * 280) to (512 * 512). On the other hand, GeoKnowledge covers 134 countries or regions, encompassing all those included in GeoExplain.
@@ 3.3 Data Analysis
Figure 2: Human Analysis on GeoExplain and Other Multimodal Reasoning Dataset
@@
4 SightSense
Figure 3: Architecture of SightSense: The first stage is Visual Clue Detection which detects identifiable visual clues. The second stage is Multimodal Knowledge Retrieval to incorporate external geographical knowledge. The final stage is to concatenate visual clues, global context and knowledge snippets as prompt to generate reasoning.
@@ 4.1 Problem Statement
$L^{*},E^{*}=\arg\max_{L,E}P(L,E|P,K;\theta)$ (3)
4.2 Model Architecture
In order to fully leverage visual information at different levels of granularity, SightSense uses a three-stage method. Figure 3 illustrates the overall workflow. The three stages are Visual Clues Detection, Multimodal Knowledge Retrieval and Reasoning Generation respectively. In the first stage, potentially identifiable local elements are extracted from street view images $\mathbf{P}$ by using an open-vocabulary object detector. The second stage retrieves relevant geographical knowledge $KI,S$ based on these visual clues. Finally, the reasoning stage integrates global context, detected visual clues and retrieved external knowledge to generate informative answer and explanation L,E{L,E}.
4.3 Visual Clue Detection